OpenStudy (priscillalynnx3):

Choose the slope-intercept equation of the line that passes through the point (-2, 4) and is parallel to y = -3x + 6.

OpenStudy (e.mccormick):

Because it is parallel, what does that tell you about the slopes of the two lines?

OpenStudy (e.mccormick):

Well, they have th slope of one and they are being asked about another. But I am trying to find where the problem is.

OpenStudy (priscillalynnx3):

it asks me for the equation of that point, and that is parallel to that line.

OpenStudy (e.mccormick):

Yes, and I am asking do you know what parallel means in regards to slope?

OpenStudy (e.mccormick):

Id this was the graph of two parallel lines, any random two lines, would you say the slope is the same, different, or what? 

OpenStudy (priscillalynnx3):

yes., the slopes would be the same for parallel lines.

OpenStudy (e.mccormick):

Great! So, you have the equation of the first line. That has the slope. Do you know what the slope is there?

OpenStudy (e.mccormick):

Once you know the slope, do you know the point-slope formula?

OpenStudy (priscillalynnx3):

y – y1 = m(x – x1)

OpenStudy (e.mccormick):

Perfect. So, what pat of the question go where in that, or is that the problem?

OpenStudy (priscillalynnx3):

i just cant find the equation

OpenStudy (e.mccormick):

The key information is in: point (-2, 4) and is parallel to y = -3x + 6. Now, do you know what the slope of the first line is?

OpenStudy (e.mccormick):

\(y=mx+b\) So what would the slope be?